OpenStudy (anonymous):

How would you set up the y=mx+b with (0, -5) and (3, -4)?

OpenStudy (anonymous):

y1-y2=m (x1-x2) so... -5+4=m(0-3) -1=m(-3) m=1/3

OpenStudy (anonymous):

What about b?

OpenStudy (anonymous):

b is the (y) intercept, so you setup the variable (y) to 0 into the original equation

OpenStudy (anonymous):

Can you show me? Sorry, I don't remember my algebra all that much.

OpenStudy (anonymous):

what was your original equation? or they are asking just for the slope?

OpenStudy (anonymous):

They're asking for the slope, but can you find the b with that pair?

OpenStudy (anonymous):

yes :) it was given to you!!! remember (0, -5) so the equation would be y=(1/3)x-5 because m=1/3 and b=-5 good luck ;) !!!!